Please help.
How many pounds of hamburger that costs $1.60 per pound must be mixed with 70 pounds of hamburger that costs $2.10 per pound to make a mixture that costs $1.70 per pound.

Answer by KnightOwlTutor(293)

X=lbs of hamburger that costs$1.60
70=lbs of hamburger that costs $2.10
Even thought the cost varies the total quantity of meat is constant.
1.60x+(2.10)(70)=1.70(x+70)
1.60x+147=1.70x+119
Subtract 1.60x from both sides
subtract 119 from both sides
28=.1x
divide both sides by .1
280=x

Let's doublecheck our answer by replacing the x with 280
1.6(280)+147=(350)1.7
448+147=595
595=595